OpenStudy (anonymous):

Which answer best describes the solution to the equation 5z + 4 = 4 + 5z? A. one solution B. no solution C. an infinite number of solutions D. z = 20

OpenStudy (igreen):

Subtract 5z to both sides..what's 5z - 5z?

OpenStudy (anonymous):

0 right?

OpenStudy (uri):

They get cancelled.

OpenStudy (anonymous):

So i would say B?

OpenStudy (igreen):

Yes..well your equation is 5z + 4 = 4 + 5z It's the same thing on both sides..therefore they're the same line..so infinitely many solutions.

OpenStudy (igreen):

5z + 4 = 4 + 5z They're both 5z + 4.

OpenStudy (anonymous):

Okay so...C then

OpenStudy (igreen):

Yep.

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

Thanks